a rectangle has a length of 4 feet and an area of 36 feet square what is the width of this rectangle
Found 2 solutions by drj, rfer:
Answer by drj(1380)

A rectangle has a length of 4 feet and an area of 36 feet square what is the width of this rectangle.

Step 1. Area = length * width or A=L*W where A is the area, L is the length, and W is the width

Step 2. Then W=A/L=36/4= 9 feet.

Step 3. ANSWER: The width is 9 feet in length.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J

Answer by rfer(16322)

A=lw
w=A/l
w=36/4
w=9